The average travel time from your residence to your school is 35 minutes with a standard deviation of 10 minutes. If you want to be 99% certain that you will not be late for your first 8:00 am class, what is the latest time you should leave home? Assume that travel time is normally distributed.

1
Expert's answer
2021-11-07T18:00:22-0500

Solution:

"P(Z<z)=0.99."

"\Rightarrow z=2.326" [Using z-score table]

"\Rightarrow \frac{x-35}{10}=2.326"

"\Rightarrow x=2.325\times10+35=58.25"

The latest time you should leave home: "8:00-58.25=7:01.75\approx7:02."